OpenStudy (anonymous):

log 16 x =3/2 solve for x logarithms

OpenStudy (mertsj):

\[\log_{16}x=\frac{3}{2} \]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (johnweldon1993):

I believe that is correct @Mertsj So the question is basically asking..... \[16^{3/2} = x\] So what is 16 to the 3/2 power? Well that is the same thing as saying \[\sqrt{16}^3\] So what is the square root of 16.....then what is THAT number to the 3rd power? This will be your 'x'